jeff ate 2/3 of a pizza if he only started with 3/4 what fraction of the pizza is left
Answer by JulietG(1812)

The first thing we have to do is to find the greatest common denominator of 3 and 4. It's 12 (I just multiplied 3 * 4).
Now find the equivalent.
2/3 = how many 12ths? Just multiply the top and bottom by 4 to get 8/12
3/4 = how many 12ths? Multiply the top and bottom by 3 to get 9/12
Now it becomes simple subtraction.
9/12 - 8/12 = 1/12
Conclusion: Jeff is a glutton.